OpenStudy (anonymous):

Evaluate the function for the given value of x. f(x) = -6x2 - 4x - 4 , x = - 2 A. f( - 2) = -36 B. f( - 2) = -17 C. f( - 2) = -20 D. f( - 2) = -28

OpenStudy (anonymous):

f(x) = -6x2 - 4x - 4 , x = - 2 f(-2) = -6(-2)2 - 4(-2) - 4 f(-2) = -6(4) +8 - 4 f(-2) = -24 +8 - 4 = -20

OpenStudy (anonymous):

isnt she sexy? lovin the pink

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

she is, but we are here for math not flirting

OpenStudy (anonymous):

well 1 + 1 does = 2 ;)

OpenStudy (anonymous):

bahaha good one

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):

ok what is x^2*y" + 5xy' - 10y = tan(x)

OpenStudy (anonymous):

no idea

OpenStudy (anonymous):

not a clue